OpenStudy (anonymous):

How would I do this? The diameter of a circle has endpoints P(-10, -2) and Q(4, 6). a. Find the center of the circle. b. Find the radius, and if the answer is not an integer, then express it in radical form. c. Write an equation for the circle.

OpenStudy (anonymous):

it will be a. but im not positive

OpenStudy (anonymous):



OpenStudy (anonymous):

ok then it will be b.

OpenStudy (anonymous):

do I understand this right? a.) then M=((4+(-10))/2,(6+(-2)/2) hence M=(-3,2)

OpenStudy (anonymous):

I think that is right! @WambelGy

OpenStudy (anonymous):

for b.) distance of M to Q: sqrt((4-(-3))^2+(6-2)^2)

OpenStudy (anonymous):

So it is 8.06?

OpenStudy (anonymous):

or do i not solve it and leave it as you had it above?

OpenStudy (anonymous):

maybe you should leave it like above

OpenStudy (anonymous):

I mean like I had it

OpenStudy (anonymous):

Okay, what about C?

OpenStudy (anonymous):

I was thinking of a parametric equation... but for a 'normal' equation it should be something like this: (x-(-3))^2+(y-2)^2=r^2 (=(4-(-3))^2+(6-2)^2 = 7^2+4^2 = 65)

OpenStudy (anonymous):

Wait so which one is it?

OpenStudy (anonymous):

(x-(-3))^2+(y-2)^2 = 65